# Mastering the Craft of Data Synthesis for CodeLLMs

### v2

@@ Abstract
Meng Chen1  Philip Arthur1  Qianyu Feng2  Cong Duy Vu Hoang3 Yu-Heng Hong1  Mahdi Kazemi Moghaddam1  Omid Nezami1  Thien Nguyen2 Gioacchino Tangari1  Duy Vu2  Thanh Vu1  Mark Johnson1 Krishnaram Kenthapadi1  Don Dharmasiri2  Long Duong1  Yuan-Fang Li1 🖂 🖂 yuanfang.li@oracle.com 1 Oracle Health and AI (OHAI), Oracle 2 Oracle Cloud Infrastructure (OCI), Oracle 3 Oracle Analytics Cloud (OAC), Oracle
@@ 1 Introduction
While generating synthetic datasets for code-related tasks may appear straightforward, achieving both high accuracy and sufficient diversity is a complex process requiring meticulous design and advanced techniques Gandhi et al. (2024). This makes a systematic exploration of LLM-driven synthetic data generation both essential and timely. Although there are survey papers in the fields of general data engineering Liu et al. (2024b); Long et al. (2024); Wang et al. (2024c); Ding et al. (2024a) and code intelligence Wan et al. (2023); Jiang et al. (2024); Zhang et al. (2024d); Sun et al. (2024a), there is a notable gap in literature focusing specifically on data synthesis and filtering techniques for code-related tasks. To fill this gap, we present a targeted review of recent advancements in synthetic data generation and filtering for training CodeLLMs, covering over 50 recent works across 23 topic categories from the past two years. The techniques discussed are organized into a taxonomy (Fig. 1) and analyzed in terms of their motivation, methodologies, and key contributions. Our goal is to provide an in-depth overview of the current state of the field, highlight key challenges, and offer insights to guide researchers and practitioners in building efficient and robust CodeLLMs through effective data engineering practices.
The remainder of this paper is structured as follows. In Section 2, we provide an overview of the data curation pipeline, define the scope and taxonomy of this survey, and highlight key distinctions from other related works. Sections 3 and 4 cover the core techniques for data synthesis and filtering, respectively. In Section 5, we discuss key challenges and outline potential future research directions. We present our conclusions in Section 6. Additionally in Appendix A, we offer a practical guide with best practices and considerations for selecting techniques and models.
@@ 2 Preliminaries and Related Works
In this section, we outline the data curation pipeline, define the survey’s scope and taxonomy, and highlight key differences from recent surveys.
@@ 2.1 The Data Curation Pipeline
Data curation, which aims to ensure datasets are of high quality, diverse, relevant, and available, is crucial to the success of CodeLLMs. The data curation process typically involves four key steps (cf. Figure 2 in Appendix A). (1) Seed Input Collection: Before synthesizing data, a small set of seed samples (e.g. problem-solution pairs), unlabeled inputs (e.g. code snippets), or human-written instructions (e.g. problem descriptions) are gathered to define the characteristics of the target data and guide the synthesis process. (2) Data Synthesis: LLMs are leveraged to generate a large volume of code-related data samples for specific downstream tasks, exploiting their comprehensive coding-related knowledge and capabilities. (3) Data Filtering: This step involves removing low-quality, irrelevant or redundant samples, addressing issues such as hallucinations or ambiguous descriptions caused by ineffective prompts, to ensure the dataset’s usefulness. (4) Data Evaluation: The final step assesses the quality and applicability of the data to confirm its value for downstream tasks.
2.2 Survey Scope and Taxonomy
@@ 6 Conclusion
Due to page limits, we may not have included all relevant works and technical details. The primary studies we included are mostly 2022 onwards. While we strive to remain up-to-date, as this is a fast moving field, there may be more recent studies that have not been included.
@@
Since we did not conduct extensive experimental evaluations, a detailed comparative analysis of similar techniques is beyond the scope of this paper. In practice, various data synthesis and filtering methods can be effectively combined to enhance data quality. Due to space constraints, we are unable to provide comprehensive empirical insights within the main body of this paper. For more detailed practical guidance, please refer to Appendix A.
Appendix A Practical Guidance
In this section, we outline a comprehensive pipeline for synthetic data generation tailored to CodeLLMs, offering practical guidance for researchers entering this domain. Building on the discussion in Section 2.1, we detail the pipeline stages in accordance with the data curation life cycle, as depicted in Figure 2. Additionally, we recommend several cost-effective yet high-quality large language models (LLMs) for the code-related data generation.
Figure 2: Practical guidance for the code related data generation pipeline, which encompasses: (Step 1) seed data collection, (Step 2) data synthesis, (Step 3) data filtering, and (Step 4) data evaluation. We also recommend several large language models (LLMs) that offer strong performance while maintaining a balanced cost.
A.1 Seed Data Collection
The initial step in synthetic data generation is to gather seed data, which can be either labeled (e.g. problem-solution pairs) or unlabeled (e.g. code snippets, API documentation). For instance, when using fine-tuning methods, seed data can be categorized into three types: (1) Instructions, which describe the requirements of the code-related tasks. For example, in the context of code generation, an instruction might be “write a Python program that generates a random password of 8 characters”. These instructional seeds can be manually crafted through crowd-sourcing efforts. It is recommended that the instructional seed data covers a broad range of tasks, with each task offering 1 or 2 variants to enhance diversity. (2) Code snippets, typically gathered from open-source platforms such as GitHub, based on the programming languages relevant to the task. If licensing permits, proprietary codebases can also serve as seed data, but sensitive information, such as personal names, phone numbers, addresses, or financial details, must be anonymized. If code in the target language is unavailable, snippets from similar languages may be utilized as a substitute. (3) Documentation, which is particularly useful for low-resource programming languages where both introductory materials and human-written code repositories may be scarce. In these cases, online API documentation can be leveraged as seed data, as it often contains valuable information on programming syntax and examples akin to textbook content.
A.2 Data Synthesis
The choice of data synthesis techniques depends on the type of seed data being utilized.
For instruction-only seed data, the initial step is to expand the instruction set into more natural, fluent, and diverse variants. Techniques such as Self-Instruct Chaudhary (2023), WizardCoder Luo et al. (2024), Auto Evol-Instruct Zeng et al. (2024), and AIEV-INSTRUCT Lei et al. (2024) are recommended as starting points. Once a sufficient number of instruction variants have been generated, LLMs can be prompted with each instruction to generate corresponding responses or solutions, forming instruction-solution pairs.
For seed data consisting solely of code snippets, Magicoder Wei et al. (2024) can be employed to stimulate LLMs to generate problem-solution pairs simultaneously. Notably, Magicoder has the capability to generate code in languages different from those found in the seed data.
When working with documentation-based seed data, approaches like those in Cheng et al. (2024) can be used to extract question-answer pairs through instruction-synthesizers, which interpret the underlying knowledge in the documentation and convert it into a fine-tuning-friendly format.
It is important to note that the aforementioned techniques are typically applied in data synthesis pipelines for supervised fine-tuning. Additionally, by using both strong and weak LLMs, researchers can generate a range of solutions, including both high- and low-quality responses, to construct datasets suitable for preference alignment training (PAT) Weyssow et al. (2024).
A.3 Data Filtering
Once the raw synthetic data is generated, it is essential to perform iterative rounds of data filtering to enhance both the quality and diversity of the dataset. The following steps are recommended: (1) A combination of filtering techniques can be employed to assess the quality of synthetic data samples. LLM-based filtering methods, such as those in Dubey et al. (2024), can be initially utilized to predict a quality score for each sample. Execution-based filtering techniques Liu et al. (2024a) should then be applied to obtain feedback from interpreters. (2) Based on the quality scores and execution feedback, LLM-based code correction Wadhwa et al. (2023) can be incorporated to refine the synthetic code and retain more usable data. (3) Following this step, filtering based on small models Li et al. (2024) and rules Li et al. (2023a) can be used to eliminate poor-quality data and efficiently de-duplicate the dataset. (4) Lastly, rigorous data decontamination should be conducted, both at the surface level Li et al. (2023a) and semantic level Riddell et al. (2024), to ensure the integrity of the model evaluation process.
A.4 Data Evaluation
Evaluating the effectiveness of synthetic data for training code-focused LLMs typically involves experimenting with various dataset combinations. Synthetic datasets can be categorized based on multiple dimensions, such as domain specificity, seed data sources, or the teacher LLMs used during generation. By conducting comprehensive ablation studies and data combination experiments, the contribution of each dataset to the overall performance can be assessed, facilitating the selection of the most effective datasets for model training. Moreover, the challenge of data mixing is crucial. In addition to traditional heuristics or manually assigned weights, recent advances in offline data mixing Xie et al. (2023) and online data mixing Albalak et al. (2023) methods offer promising alternatives. These approaches can be applied at different stages of CodeLLM development, including pre-training, supervised fine-tuning, and preference alignment training (PAT).
A.5 LLM Selection Considerations
@@
Given the central role of LLMs play in both the synthesis and filtering processes, selecting the appropriate LLM is crucial for optimizing performance. Based on our empirical observations, we recommend several commonly used models. If cost is not a limiting factor, GPT-4 or GPT-4o [^1] https://platform.openai.com/docs/models/gpt-4o remains a top-tier choice, consistently delivering high-quality results. Alternatively, the open-sourced Llama 3.1-405B model [^2] https://huggingface.co/meta-llama/Llama-3.1-405B, particularly its int4 quantized version [^3] https://huggingface.co/neuralmagic/Meta-Llama-3.1-405B-Instruct-quantized.w4a16, is a strong candidate. It can be efficiently run on four H100 GPUs, offering a good trade-off between quality and computational cost. Another viable option is the recently released Qwen2.5-72B-Instruct model [^4] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, known for its fast execution and proficiency in code-related tasks. Lastly, DeepSeek-Coder-V2-Instruct [^5] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Instruct, an open-source Mixture-of-Experts (MoE) CodeLLM, is also worth considering. It has demonstrated comparable performance to GPT-4-Turbo in code-specific tasks.
@@
Besides, researchers and practitioners should be aware of the specific synthetic data usage policies associated with various LLMs, particularly when building commercial products. For instance, Llama 2 and 3 restrict the use of generated outputs for training other AI models, whereas Llama 3.1 and 3.2 have updated these policies [^6] https://www.llama.com/faq/. The Qwen model, on the other hand, requires explicit attribution, such as prominently displaying ‘Built with Qwen’ or ‘Improved using Qwen’ in product documentation when its outputs are used for creating, training, fine-tuning, or improving an AI model [^7] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct/blob/main/LICENSE.
@@
For inference, vLLM [^8] https://github.com/vllm-project/vllm is a fast and user-friendly library that supports high-throughput serving with various decoding algorithms, such as parallel sampling and beam search. It is compatible with most popular open-source models on Hugging Face, including Transformer-based LLMs (e.g. Llama Dubey et al. (2024)), Mixture-of-Experts LLMs (e.g. Mixtral [^9] https://huggingface.co/mistralai), Embedding Models (e.g. E5-Mistral [^10] https://huggingface.co/intfloat/e5-mistral-7b-instruct), and Multi-modal LLMs (e.g. LLaVA Liu et al. (2023a)). It also allows for offline batched inference on datasets or sending requests through an OpenAI-compatible API server, making it highly convenient for large-scale data experimentation.

### v3

@@ Abstract
Meng Chen🖂  Philip Arthur  Qianyu Feng  Cong Duy Vu Hoang Yu-Heng Hong  Mahdi Kazemi Moghaddam  Omid Nezami  Thien Nguyen Gioacchino Tangari  Duy Vu  Thanh Vu  Mark Johnson Krishnaram Kenthapadi  Don Dharmasiri  Long Duong  Yuan-Fang Li🖂 🖂 {meng.c.chen, yuanfang.li}@oracle.com Oracle Corporation
@@ 1 Introduction
Figure 1: Practical guidance for the code related data generation pipeline. We also recommend several large language models (LLMs) that offer strong performance while maintaining a balanced cost (Appendix A.2).
While generating synthetic datasets for code-related tasks may appear straightforward, achieving both high accuracy and sufficient diversity is a complex process requiring meticulous design and advanced techniques Gandhi et al. (2024). This makes a systematic exploration of LLM-driven synthetic data generation both essential and timely. Although there are survey papers in the fields of general data engineering Liu et al. (2024b); Long et al. (2024); Wang et al. (2024c); Ding et al. (2024a) and code intelligence Wan et al. (2023); Jiang et al. (2024); Zhang et al. (2024d); Sun et al. (2024a), there is a notable gap in literature focusing specifically on data synthesis and filtering techniques for code-related tasks. To fill this gap, we present a targeted review of recent advancements in synthetic data generation and filtering for training CodeLLMs, covering over 50 recent works across 23 topic categories from the past two years. The techniques discussed are organized into a taxonomy (Fig. 2) and analyzed in terms of their motivation, methodologies, and key contributions. We also maintain a GitHub repository [^1] https://github.com/chenmengdx/awesome-data-synthesis-for-code-llm to collect open-source synthesis datasets for CodeLLMs and track recent advancements. Our goal is to provide an in-depth overview of the current state of the field, highlight key challenges, and offer insights to guide researchers and practitioners in building efficient and robust CodeLLMs through effective data engineering practices.
@@ 2.1 The Data Curation Pipeline
Data curation, which aims to ensure datasets are of high quality, diverse, relevant, and available, is crucial to the success of CodeLLMs. The data curation process typically involves four key steps (cf. Figure 1 in  6). (1) Seed Input Collection: Before synthesizing data, a small set of seed samples (e.g. problem-solution pairs), unlabeled inputs (e.g. code snippets), or human-written instructions (e.g. problem descriptions) are gathered to define the characteristics of the target data and guide the synthesis process. (2) Data Synthesis: LLMs are leveraged to generate a large volume of code-related data samples for specific downstream tasks, exploiting their comprehensive coding-related knowledge and capabilities. (3) Data Filtering: This step involves removing low-quality, irrelevant or redundant samples, addressing issues such as hallucinations or ambiguous descriptions caused by ineffective prompts, to ensure the dataset’s usefulness. (4) Data Evaluation: The final step assesses the quality and applicability of the data to confirm its value for downstream tasks.
@@ 3.3 Specific Tasks
Data Synthesis Takeaways Quality & Efficiency: CodeLLMs rely on human data (e.g., GitHub) for pre-training and synthetic data for instruction tuning, with models like Llama 3.1 and Qwen2.5-Coder favoring the latter for its efficiency. Key Enhancements: Improving synthetic data via interpreter feedback, better seed selection, and reasoning steps enhances CodeLLMs. Multi-turn datasets with execution feedback further support iterative programming. Future work explores agent-like learning. Task Adaptation: Synthetic data effectively tailors CodeLLMs to specific tasks, though challenges remain in supporting low-resource languages and version-specific code generation.
@@ 4.5 Decontamination
Data Filtering Takeaways Optimized Filtering: A hybrid of rule-based and model-based techniques balances computational efficiency and dataset size. Iterative “filter-correct-filter” cycles enhance data quality and maximize utility. Dataset Composition: Beyond filtering, strategically mixing datasets in optimal ratios improves diverse capabilities, including reasoning, mathematical proficiency, and general language skills in CodeLLMs. Decontamination for Robust Evaluation: Ensuring unbiased CodeLLM evaluation requires rigorous decontamination. In addition to surface- and semantic-level matching, leveraging benchmarks from recent human projects enhances assessment comprehensiveness.
@@ 5 Challenges and Future Directions
6 Practical Guidance
This section introduces a streamlined pipeline (Fig. 1) for synthetic data generation in CodeLLMs, providing practical guidance for researchers.
6.1 Seed Data Collection
The first step in synthetic data generation is collecting seed data, which can be labeled (e.g., problem-solution pairs) or unlabeled (e.g., code snippets, API documentation). For fine-tuning, seed data falls into three categories: (1) Instructions, which define code-related task requirements. For instance, a code generation instruction might be “write a Python program that generates a random password of 8 characters”. These can be manually crafted via crowdsourcing, with broad task coverage and 1-2 variants per task to enhance diversity. (2) Code snippets, sourced from open platforms like GitHub, based on relevant programming languages. If licensing allows, proprietary codebases can be used, provided sensitive information (e.g., personal names, contact details) is anonymized. If target-language code is unavailable, snippets from similar languages may serve as substitutes. (3) Documentation, particularly valuable for low-resource languages lacking introductory materials and scarce human-written repositories. Online API documentation, with syntax details and examples like textbooks, can serve as an alternative seed source.
6.2 Data Synthesis
The choice of data synthesis techniques depends on the seed data type. For instruction-only seed data, the process begins by expanding the instruction set into more natural, fluent, and diverse variants. Techniques such as Self-Instruct Chaudhary (2023), WizardCoder Luo et al. (2024), Auto Evol-Instruct Zeng et al. (2024), and AIEV-INSTRUCT Lei et al. (2024) serve as effective starting points. Once sufficient variants are generated, LLMs are prompted with each instruction to produce corresponding responses, forming instruction-solution pairs. For seed data comprising only code snippets, Magicoder Wei et al. (2024) facilitates the simultaneous generation of problem-solution pairs and can produce code in languages beyond those present in the seed data. For documentation-based seed data, methods like Cheng et al. (2024) leverage instruction-synthesizers to extract question-answer pairs by interpreting the underlying knowledge and formatting it for fine-tuning. These techniques are typically employed in supervised fine-tuning pipelines. Besides, leveraging both strong and weak LLMs generates diverse responses, including high- and low-quality solutions, supporting preference alignment training Weyssow et al. (2024).
6.3 Data Filtering
After generating raw synthetic data, iterative filtering is crucial to enhance dataset quality and diversity. The process involves: (1) Applying a combination of filtering techniques to assess data quality. LLM-based methods Dubey et al. (2024) initially predict quality scores, followed by execution-based filtering Liu et al. (2024a) to gather interpreter feedback. (2) Leveraging LLM-based code correction Wadhwa et al. (2023) to refine synthetic code based on quality scores and execution feedback. (3) Employing small model-based Li et al. (2024) and rule-based Li et al. (2023a) filtering to remove low-quality data and eliminate duplicates efficiently. (4) Conducting rigorous data decontamination at surface Li et al. (2023a) and semantic Riddell et al. (2024) levels to ensure model evaluation integrity.
6.4 Data Evaluation
Evaluating synthetic data for training code-focused LLMs involves experimenting with diverse dataset combinations. Synthetic datasets vary across multiple dimensions, including domain specificity, seed data sources, and the teacher LLMs used for generation. Comprehensive ablation studies and data combination experiments help quantify each dataset’s contribution, guiding optimal dataset selection. A key challenge is data mixing. Beyond traditional heuristics or manually assigned weights, recent advances in offline Xie et al. (2023) and online Albalak et al. (2023) data mixing offer promising alternatives. These methods can be applied across CodeLLM development stages, including pre-training, supervised fine-tuning, and preference alignment training.
@@ 7 Conclusion
Due to page limits, we may not have included all relevant works and technical details. The primary studies we included are mostly 2022 onwards. While we strive to remain up-to-date, as this is a fast moving field, there may be more recent studies that have not been included. For the latest updates, please refer to our GitHub repository1.
@@
Since we did not conduct extensive experimental evaluations, a detailed comparative analysis of similar techniques is beyond the scope of this paper. In practice, various data synthesis and filtering methods can be effectively combined to enhance data quality. Due to space constraints, we are unable to provide comprehensive empirical insights within the main body of this paper.
Appendix A Appendix
This section provides two supplementary discussions. First, we delineate the scope of this survey, distinguishing it from related concepts such as knowledge distillation and data augmentation. Second, we recommend cost-effective yet high-quality large language models (LLMs) for code-related data generation. Given the rapid evolution of LLMs, we refer readers to our GitHub repository for the latest model updates.
A.1 Survey Scope
A.2 LLM Selection Considerations
@@
LLMs play a central role in both synthesis and filtering, making model selection critical for performance optimization. Based on empirical observations, we recommend several commonly used models. If cost is not a concern, GPT-4 or GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o remains the top choice, consistently delivering high-quality results. For a cost-effective alternative, the open-source Llama 3.1-405B [^3] https://huggingface.co/meta-llama/Llama-3.1-405B, particularly its int4 quantized version [^4] https://huggingface.co/neuralmagic/Meta-Llama-3.1-405B-Instruct-quantized.w4a16, offers a strong trade-off between quality and efficiency, running on four H100 GPUs. Another viable option is Qwen2.5-72B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, known for fast execution and strong performance in code-related tasks. Finally, DeepSeek-Coder-V2-Instruct [^6] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Instruct, an open-source Mixture-of-Experts (MoE) CodeLLM, achieves performance comparable to GPT-4-Turbo in code-specific tasks.
@@
Besides, researchers and practitioners should be aware of the specific synthetic data usage policies associated with various LLMs, particularly when building commercial products. For instance, Llama 2 and 3 restrict the use of generated outputs for training other AI models, whereas Llama 3.1 and 3.2 have updated these policies [^7] https://www.llama.com/faq/. The Qwen model, on the other hand, requires explicit attribution, such as prominently displaying “Built with Qwen” or “Improved using Qwen” in product documentation when its outputs are used for creating, training, fine-tuning, or improving an AI model [^8] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct/blob/main/LICENSE.
@@
For inference, vLLM [^9] https://github.com/vllm-project/vllm is a fast and user-friendly library that supports high-throughput serving with various decoding algorithms, such as parallel sampling and beam search. It is compatible with most popular open-source models on Hugging Face, including Transformer-based LLMs (e.g. Llama Dubey et al. (2024)), Mixture-of-Experts LLMs (e.g. Mixtral [^10] https://huggingface.co/mistralai), Embedding Models (e.g. E5-Mistral [^11] https://huggingface.co/intfloat/e5-mistral-7b-instruct), and Multi-modal LLMs (e.g. LLaVA Liu et al. (2023a)). It also allows for offline batched inference on datasets or sending requests through an OpenAI-compatible API server, providing a convenient solution for large-scale data experimentation and serving as a valuable tool for advancing research and development in this field.